# T-SCEND: Test-time Scalable MCTS-enhanced Diffusion Model

### v2

@@ F.1 Visualization of Maze experiments
F.2 Visualization of Sudoku experiments
@@
This section presents visualizations of the training and test Sudoku data in Fig. 11, and representative samples generated by different methods in Fig. 12. In the visuals, black numbers denote the condition, green numbers represent correct predictions, and red numbers represent wrong predictions. All visualizations are derived from a few representative samples. The comparison between the training and test sets clearly indicates that the tasks in the test set are significantly more difficult than those in the training set. When comparing the samples generated by different methods, it is evident that the originally trained model, regardless of the inference strategy, consistently underperforms compared to T-SCEND.

### v3

@@ H.1 Visualization of Maze experiments
H.2 Visualization of Sudoku experiments
@@
This section presents visualizations of the training and test Sudoku data in Fig. 12, and representative samples generated by different methods in Fig. 13. In the visuals, black numbers denote the condition, green numbers represent correct predictions, and red numbers represent wrong predictions. All visualizations are derived from a few representative samples. The comparison between the training and test sets clearly indicates that the tasks in the test set are significantly more difficult than those in the training set. When comparing the samples generated by different methods, it is evident that the originally trained model, regardless of the inference strategy, consistently underperforms compared to VFScale.